OpenStudy (anonymous):

4(2x − 1) − 3 = 9(x − 4)

OpenStudy (anonymous):

Simply use the distributive property then solve for x: 8x - 4 - 3 = 9x - 36 -4 + 36 -3 = 9x - 8x 29 = x Therefore, the solution is x = 29.